# Rhythmic sharing: A bio-inspired paradigm for zero-shot adaptive learning in neural networks

### v4

@@ 2.1 Unsupervised state estimation with rhythmic sharing
Figure 2: Rapid sensing of dynamic state changes A. A simulated trajectory of the 3D Thomas system (whose state variables are $x,y,z$) in orange, with the value of its internal parameter $\lambda_{\text{Thomas}}$ that was used to simulate the trajectory shown in purple. After an initial transient, $R$ remains near an equilibrium value of 1 for the entire chaotic trajectory. When $\lambda_{\text{Thomas}}$ is abruptly switched to $0.29$, the trajectory switches from a chaotic state to a periodic orbit. $R$ senses this switch immediately and settles to a new equilibrium value. B. The same procedure conducted for the 3D Lorenz system. $R$, although noisier than the Thomas system, evolves to a different equilibrium value at the exact time $\lambda_{\text{Lorenz}}$ was switched, falling to a different value as the system transitions to a dead signal. C. $R$ for the Mackey-Glass system, whose internal parameter $\lambda_{\text{Mackey}}$ was changed sinusoidally during simulation.
@@ 2.1.1 Detecting simple state changes
We emphasize two observations about $R(t)$. First, $R$ converges to some equilibrium value whenever the input trajectory belongs to a stationary state. For instance, $R$ evolves from its initial condition 00 and reaches a steady state value $1$ within roughly two ‘cycles’ of the chaotic state ($\lambda_{\text{Thomas}}=0.18$) and maintains this value hereafter. This means that the links detected an invariant measure underlying the trajectory, even when the trajectory is chaotic; in other words, the links detected that a stationary probability distribution underlies the time evolution of the trajectory. Second, each state gives rise to a unique equilibrium value of $R$. Therefore, the evolution of $R$ mimics the evolution of $\lambda_{\text{Thomas}}$.
@@ 4.1.2 Mackey-Glass system
Here, the exact time dependence of the state variable $x$ is provided for clarity. $\tau$ is the time delay parameter responsible for inducing bifurcations of the attractor (parameterized with phase space coordinates $(x(t),x(t-\tau)$), and is recast as the parameter $\lambda_{\text{Mackey}}$ in the main text. We fix $\beta=0.2,\gamma=0.1$, and $n=10$, while we oscillate $\tau$ sinusoidally between $\tau=20$ and $\tau=24$ with a slow angular frequency ($=0.007$) during the simulation of our data.
@@ 4.1.3 Lorenz system
Here, the parameter $\rho$ is responsible for bifurcations of the system and is recast as the parameter $\lambda_{\text{Lorenz}}$ in the main text. We fix the other parameters to values $\sigma=10$, $\beta=8/3$ for the duration of the entire simulation, while we switch the value of $\rho$ abruptly as in the case for the Thomas system. For $\rho\gtrapprox 24.06$, the system permits a globally stable strange attractor that coexists with two stable fixed points, but the former’s trajectories never hit the latter’s basin boundary. For $13.93\gtrapprox\rho\lessapprox 24.06$, the strange attractor turns into a chaotic saddle, which also coexists with the fixed points, but trajectories of the saddle will hit the saddle boundary after some probabilistic time[16].
@@ 4.2 Rhythmic sharing
where $\circ$ denotes the Hadamard product. $\bm{\Phi}$ is a vector of phases $[\phi_{1}\,...\,\phi_{N_{l}}]^{\text{T}}$ with $N_{l}$ being the total number of links. $\bm{\omega}_{0}$ is a vector of initially imposed natural frequencies of the links. The vector of local mean fields $\bm{\Psi}(t)\in\mathbb{R}^{N_{l}}$ that each link is coupled to is defined by
@@
where the activation function $\tanh[\cdot]$ is applied to each vector equation component in its argument separately. It is to be noted that when we refer to a node, neuronal, or reservoir state, we mean a point of the vector $\mathbf{n}\in(-1,1)^{N_{n}}$; all other uses of the term ‘state’ specifically refers to an invariant set, such as an attractor. $\alpha$ is the leakage hyperparameter that sets the rate of update and memory of $\mathbf{n}$, $\mathbf{A_{n}}\in\mathbb{R}^{N_{n}\times N_{n}}$ is another random adjacency matrix between the nodes (scaled by its spectral radius, a hyperparameter), $\mathbf{W}_{\text{in}}\in\mathbb{R}^{\text{N}_{n}\times\text{N}_{u}}$ is the random input matrix whose elements are chosen uniformly from the closed interval $[-\delta,\delta]$ where $\delta$ is the input scaling hyperparameter, $\mathbf{u}(t)\in\text{N}_{u}$ is the input, $\bm{\xi}\in\mathbb{R}^{N_{n}}$ is a node bias hyperparameter that can be used to shift the activation function of the neuronal nodes. While eq. 6 subsumes the nonlinearity of activation to each node, we note that it is equivalent to other recurrent neural network update equations up to a coordinate transformation[23]. As such, we do not lose generality by using an echo state-type network.
@@
$\begin{split}\bm{\underline{\Phi}}(t)_{ij}:=\left\{\begin{matrix}\underline{\phi}_{j\rightarrow i}&:\,\mathbf{A}_{{\mathbf{n}}_{ij}}\neq 0\\ 0&:\,\mathbf{A}_{{\mathbf{n}}_{ij}}=0\end{matrix}\right.\end{split}$ (9)
@@ 4.2.1 Oscillation death for isolated links
[…]
@@
For $|\omega_{0}|<|\varepsilon_{0}|$, $\Delta$ and $\xi_{0}$ are purely imaginary, making both the amplitude and the argument of the nested $\tan$ in the expression for $\phi(t)$ also purely imaginary. As such, $\phi(t)$ evolves as $2\tan^{-1}[a+b\tanh(ct+d)]$, where $a,\,b,\,c,\,d$ are real constants. Since $\tanh(ct+d)$ possesses a real limit for large $t$, $\phi(t)$ converges to a real limit, i.e.,the oscillation has quenched.
On the other hand, consider $|\omega_{0}|>|\varepsilon_{0}|$. Then $\Delta$ and $\xi_{0}$ are purely real, and $\phi(t)$ evolves as $2\tan^{-1}[a+b\tan(ct+d)]$, where $a,\,b,\,c,\,d$ are reused to denote arbitrary real constants. Because the argument does not have a real limit for large $t$, the phase $\phi(t)$ also does not. Therefore, $|\omega_{0}|<|\varepsilon|$ for oscillation death to occur in isolated links.
@@ 4.3.1 Training
$\begin{split}\operatorname*{arg\,min}_{\mathbf{W}_{\text{out}}}\bigg{[}\sum_{t=t_{\text{warm}}+\Delta t}^{t_{\text{warm}}+\tau_{\text{train}}\Delta t}||\mathbf{W}_{\text{out}}\mathbf{n}(t)-\mathbf{u}_{\text{train}}(t)||^{2}+\beta\,\text{tr}(\mathbf{W}_{\text{out}}\mathbf{W}_{\text{out}}^{\text{T}})\bigg{]}\end{split}$ (13)
@@ 4.3.3 State targeting and prediction by steering $(R,\langle\Phi\rangle)$
for $\tau>2\pi/\Omega$ to ensure the space of static link strength configurations given $R_{0}$ is fully sampled during this process, and where $h$ is an error function of choice. The warm-up duration of the prediction is chosen such that all the targeting of the mean-field variables aforementioned happens before the end of the warm-up period, i.e.,$t_{R_{0}}+\tau\leq\tau_{\text{warm}}\Delta t$. For most systems, as we have used, a natural candidate is the root mean square error: $h^{2}=||\mathbf{\hat{W}}_{\text{out}}\mathbf{n}(\bm{\Phi},t)-\mathbf{u}_{\text{test}}(t)||^{2}$, but if the topology of the state does not change considerably with the system’s bifurcation parameter or the system is contaminated with noise, an error function that can more finely resolve differences in states should be formulated and used. In summary,

### v5

@@ 2.1 Unsupervised state estimation with rhythmic sharing
Figure 2: Rapid sensing of dynamic state changes A. A simulated trajectory of the 3D Thomas system (whose state variables are $x,y,z$) in orange, with the value of its internal parameter $\lambda_{\text{Thomas}}$ that was used to simulate the trajectory shown in purple. After an initial transient, $R$ remains near an equilibrium value of 1 for the entire chaotic trajectory. When $\lambda_{\text{Thomas}}$ is abruptly switched to $0.29$, the trajectory switches from a chaotic state to a periodic orbit. $R$ senses this switch immediately and settles to a new equilibrium value. B. The same procedure conducted for the Mackey-Glass system, whose internal parameter $\lambda_{\text{Mackey}}$ was changed sinusoidally during simulation. C. (Left) $R$ for the 3D Lorenz system. $R$, although noisier than the Thomas system, evolves to a different equilibrium value at the exact time $\lambda_{\text{Lorenz}}$ was switched, falling to a different value as the system transitions to a dead signal. (Right) The power spectral densities (PSD) of the time series of each state variable, corresponding to when the system is at stable chaos (green) and unstable, or transient, chaos (orange). That their spectral differences are also not obvious contributes to the difficulty in sensing this state transition.
@@ 2.1.1 Detecting simple state changes
We emphasize two observations about $R(t)$. First, $R$ converges to some equilibrium value whenever the input trajectory belongs to a stationary state. For instance, $R$ evolves from its initial condition 0 and reaches a steady state value $1$ within roughly two ‘cycles’ of the chaotic state ($\lambda_{\text{Thomas}}=0.18$) and maintains this value hereafter. This means that the links detected an invariant measure underlying the trajectory, even when the trajectory is chaotic; in other words, the links detected that a stationary probability distribution underlies the time evolution of the trajectory. Second, each state gives rise to a unique equilibrium value of $R$. Therefore, the evolution of $R$ mimics the evolution of $\lambda_{\text{Thomas}}$.
@@ 4.1.2 Mackey-Glass system
Here, the exact time dependence of the state variable $x$ is provided for clarity. $\tau$ is the time delay parameter responsible for inducing bifurcations of the attractor (parameterized with phase space coordinates $(x(t),x(t-\tau)$), and is recast as the parameter $\lambda_{\text{Mackey}}$ in the main text. We fix $\beta=0.2,\gamma=0.1$, and $n=10$, while we oscillate $\tau$ sinusoidally between $\tau=20$ and $\tau=24$ with a slow angular frequency (=0.007=0.007) during the simulation of our data.
@@ 4.1.3 Lorenz system
Here, the parameter $\rho$ is responsible for bifurcations of the system and is recast as the parameter $\lambda_{\text{Lorenz}}$ in the main text. We fix the other parameters to values $\sigma=10$, $\beta=8/3$ for the duration of the entire simulation, while we switch the value of $\rho$ abruptly as in the case for the Thomas system. For $\rho\gtrapprox 24.06$, the system permits a globally stable strange attractor that coexists with two stable fixed points, but the former’s trajectories never hit the latter’s basin boundary. For $13.93\lessapprox\rho\lessapprox 24.06$, the strange attractor turns into a chaotic saddle, which also coexists with the fixed points, but trajectories of the saddle will hit the saddle boundary after some probabilistic time[16].
@@ 4.2 Rhythmic sharing
where $\circ$ denotes the Hadamard product. $\bm{\Phi}$ is a vector of phases $\phi_{1}\,...\,\phi_{N_{l}}]^{\text{T}}$ with $N_{l}$ being the total number of links. $\bm{\omega}_{0}$ is a vector of initially imposed natural frequencies of the links. The vector of local mean fields $\bm{\Psi}(t)\in\mathbb{R}^{N_{l}}$ that each link is coupled to is defined by
@@
where the activation function $\tanh[\cdot]$ is applied to each vector equation component in its argument separately. It is to be noted that when we refer to a node, neuronal, or reservoir state, we mean a point of the vector $\mathbf{n}\in(-1,1)^{N_{n}}$; all other uses of the term ‘state’ specifically refers to an invariant set, such as an attractor. $\alpha$ is the leakage hyperparameter that sets the rate of update and memory of $\mathbf{n}$, $\mathbf{A_{n}}\in\mathbb{R}^{N_{n}\times N_{n}}$ is another random adjacency matrix between the nodes (scaled by its spectral radius, a hyperparameter), $\mathbf{W}_{\text{in}}\in\mathbb{R}^{\text{N}_{n}\times\text{N}_{u}}$ is the random input matrix whose elements are chosen uniformly from the closed interval $\delta,\delta]$ where $\delta$ is the input scaling hyperparameter, $\mathbf{u}(t)\in\text{N}_{u}$ is the input, $\bm{\xi}\in\mathbb{R}^{N_{n}}$ is a node bias hyperparameter that can be used to shift the activation function of the neuronal nodes. While eq. 6 subsumes the nonlinearity of activation to each node, we note that it is equivalent to other recurrent neural network update equations up to a coordinate transformation[23]. As such, we do not lose generality by using an echo state-type network.
@@
[…]
@@ 4.2.1 Oscillation death for isolated links
where $\Psi_{0}$ is the constant $\text{atan2}(0./0.)$, which is mathematically undefined yet computationally yields some numerical value due to floating point accuracy. $\varepsilon\equiv\varepsilon_{1}+\varepsilon_{2}(\mathbf{Q^{T}\,n}^{*})_{\phi}$ is the strength of the interaction for this particular oscillator, and which is bounded in general between $\varepsilon_{1}\pm\varepsilon_{2}$. Here, $\varepsilon$ is frozen as a constant for simplicity and to obtain an analytic solution.
@@
For $|\omega_{0}|<|\varepsilon_{0}|$, $\Delta$ and $\xi_{0}$ are purely imaginary, making both the amplitude and the argument of the nested $\tan$ in the expression for $\phi(t)$ also purely imaginary. As such, $\phi(t)$ evolves as $2\tan^{-1}[a+b\tanh(ct+d)]$, where a,b,c,da,\,b,\,c,\,d are real constants. Since $\tanh(ct+d)$ possesses a real limit for large $t$, $\phi(t)$ converges to a real limit, i.e.,the oscillation has quenched.
On the other hand, consider $|\omega_{0}|>|\varepsilon_{0}|$. Then $\Delta$ and $\xi_{0}$ are purely real, and $\phi(t)$ evolves as $2\tan^{-1}[a+b\tan(ct+d)]$, where a,b,c,da,\,b,\,c,\,d are reused to denote arbitrary real constants. Because the argument does not have a real limit for large $t$, the phase $\phi(t)$ also does not. Therefore, $|\omega_{0}|<|\varepsilon|$ for oscillation death to occur in isolated links.
@@ 4.3.1 Training
[…]
@@ 4.3.3 State targeting and prediction by steering $(R,\langle\Phi\rangle)$
[…]